Two Angles are Supplemental.One is 40 degrees more than 3 times the other. What is the other?
Answer by arunraiindian(5)

let the angle bse x so the other angle will be 90-x
acc. to ques
x=3(90-x)+40
x=270-3x+40
x=310-3x
4x=310
x=310/4
x=77.5